Prepare 25.0 mL each of 50.00 μM solutions of Cl, NO3, PO43−, and SO42− from KCl, KNO3, NaH2PO4, and K2SO4. A 50.00 μM solution of F will be provided. Run chromatograms of the individual ions and the qualitative unknown.

You cannot make up the 50.0 μM solutions directly. You will have to determine reasonable stock solution concentrations which will then be diluted.
Answer #1

reasonable stock solution concentrations which will then be diluted , to make 50.00 μM solutions of Cl, NO3, PO43−, and SO42− from KCl, KNO3, NaH2PO4, and K2SO4.

we need 25 ml solution , ==> reasonable stock solution concentration : 1250 μM

( 1: 25 dilution will be needed).

===> we prepare 500 mL of stock solution :

thus , moles of interest species in solution : 1250 micro mol /L * 0.5 L = 6.25*10-4 moles

to be dissolved in Volumetric flask of 500 mL complete the volume.

  • Cl <== from KCl : 6.25*10-4 moles * 74.55 g/mol = 0.0466 g
  • NO3 : <== from KNO3 : 6.25*10-4 moles * 101.1 g/mol = 0.0632 g
  • PO43− :<== from NaH2PO4: 6.25*10-4 moles * 120 g/mol = 0.075 g
  • SO42− : <== from K2SO4 : 6.25*10-4 moles * 174.26 g/mol = 0.1089 g